3) The experimental setup and the diffraction pattern from a single slit are given below. The screen is located at 1 meter apart from the slit wheel.- Please answer the following questions: 1 meter D1 = D2= 16.25 mm w = 0.04 mm Hint: the equation for destructive interference is given by
Answer #1

d*sin(θ) = m*λ

(a)
Now, tan(θ) = (16.25 * 10^-3)/1
θ = 0.931o

Now,
0.04 * 10^-3 * sin( 0.931o) = 1 * λ
λ = 649.93 nm

(b)
If we move the screen closer, the new angle will be θ = 1.861o
So, sin(θ) will increase,
λ is directly proportional to sin(θ),
wavelength λ will increase.

There will be no effect on D1 & D2.

Correct option - (4)